OpenStudy (anonymous):

verify secˆ2y+tanˆ2y=(1−sinˆ4y)secˆ4y

OpenStudy (anonymous):

\[\secˆ2y+\tanˆ2y=(1−\sinˆ4y)\secˆ4y\]

OpenStudy (anonymous):

I got as far as. Cross multiplying between the 2 sides is not allowed. \[\frac{1+\sin ^{2}y }{\cos ^{2}y}=\frac{1-\sin ^{4}y }{\cos ^{4}y }\]